# Theme Article: Inclusive Data Experiences

### v1

@@ Abstract
To achieve our aim, we performed an empirical experiment. We developed three narrative visualization examples that employed different interactive narrative patterns tailored specifically for engagement [8]. We randomly assigned 2400 participants to one of the three narrative visualization examples and measured their engagement using VisEngage [6]. VisEngage is a self-reporting questionnaire specifically developed to measure engagement in visualization. The age groups were split into generations. A younger audience with an age of 18-27, a middle-younger audience between the ages of 28-43, an older audience of 44-60, and finally, the oldest age cohort consisting of over 60. These age groups were determined by generation boundaries.
@@
Previous work on different age cohorts in visualization usually focused on either the very old or the very young. For example, visualization research with children has investigated pedagogical approaches for visual literacy and visualization design. Visualization research into elderly audiences has examined aspects of accessibility, comprehension and perception. A lack of understanding of different age groups needs is a known and cited dilemma in visualization research [2].
In the HCI field, age-based research has found significant differences between age groups. For example, the time taken to perform input modalities of users compared to the age group found older adults to be significantly different to their younger counterparts. Even though the results were inconclusive, strong evidence pointed towards differences between older and younger age cohorts when using devices such as smartwatches [18].
@@
We conducted a crowd-sourced study using three narrative visualization examples as a stimulus to achieve our research aim of investigating if and how end-user age impacts engagement in interactive narrative visualization. To see the interactive narrative visualization example code and raw data, please see the supplementary material here.
@@
The second narrative visualization example, Design Two, used the ‘Breaking the Fourth Wall’ interactive narrative pattern. ‘Breaking the Fourth Wall’ is a term often cited in cinema and literature disciplines. In interactive narrative visualization, a direct question is asked of the audience, normally to input personal data. This creates a ‘self-story connection,’ which has been found to encourage engagement, as it includes the user within the story [4]. The narrative visualization design example that inspired this study was a finalist in the online journalism awards. Published by the ABC Australia Story Lab, the narrative visualization is titled ‘See how global warming has changed the world since your childhood.’ In our study, Design Two broke the fourth wall by asking the user to input their name. Specifically, the user was asked to ‘please enter your first name (this data will not be stored)’ so that privacy concerns were availed with the assurance that data relating to the user’s name would not be stored. Design Two opened by stating, ‘How likely are you to be murdered by a serial killer?’ The user was asked to enter their first name click submit. Once submitted, a screen appeared with ‘[name], have you ever watched a horror movie and found yourself scared? Should you be scared of serial killers?’ This is followed by a basic bar chart explaining that ’1% of homicides in the UK are about 5.’ The narrative sequentially revealed itself as the user scrolled down the screen.’ All data sources are referenced at the end of the narrative visualization. The participants that received Design Two were asked a comprehension check question: ‘how many people earned a PhD?’ See Figure 1:B for a diagram of Design Two.
@@ Results
‘Interactive’ was mentioned 7 times by the 60+ cohort. This was in contrast to the 18-27 age cohort, who mentioned ‘interactive’ or ‘interactivity’ 49 times. We analyzed the exact phrases that participants used in the 18-27 cohort relating to interactivity. Our analysis revealed that the younger audience attributed their engagement to interactivity, for example, “yes because it was interactive” or I was engaged as it was an interactive task” (both comments from 18-27 age cohort, Design One). The interactive device in Design Two, aimed to include the audience in the story and thus encourage engagement. We coded 22 instances where the 18-24 age cohort recognized that this was the aim of the interactive device in Design 2. For example, “Yes as it was interactive and by using my name felt personal” (18-27 age cohort, Design Two). In contrast, we coded 3 instances in the 60+ age cohort specifically mentioning the interactive device in Design Two. We found further evidence of perceptiveness in Design Three from the 18-27 age cohort, where we coded 18 responses that specifically mentioned the interactive device, which in this case required clicking and exploring the interface. For example, “I did feel I was engaged in the data story as I had to click to find the information as well as scroll for more information” (18-27 age cohort, Design Three). The goal of the interactive device in Design One, ‘Make a Guess’ was to illustrate a mismatch between audience expectation and reality. We coded 17 responses from the 18-27 age cohort that specifically mentioned the interactive device. For example, ‘The decline is higher’ which made me feel engaged” (18-27 age cohort, Design One). In contrast the 60+ age cohort mentioned the interactive device in Design One 6 times. While we observed that 18-24 age cohort was relatively more aware of the interactive devices this does not mean other age groups were oblivious, only their perceptive feedback was less frequent. For example, “I felt engaged as there was an interactive question where I could enter what I thought to be the answer. This made the impact of learning the true answer heavier as I was engaging with the story” (44-59 age cohort, Design One) or “I found the personality types interesting flip over to read and to associate the descriptions with the images you gave for the personality” (60+ age cohort, Design Three). When recognized by the older age cohorts, there were 17 instances where the interactive device had the opposite effect of encouraging engagement and reported as a distraction or confusion, “I was distracted by, initially, not realizing I had to scroll down the box to gain more information” (60+ age cohort, Design Two) or “I found the chart near the end distracting as it seemed confusing.” (44-59 age cohort, Design Three). In the 18-27 age cohort, there were 2 observable reports of the interactive devices causing confusion and none causing distraction. The two reports were varied on the reasoning for why the interactive narrative visualization was deemed confusing. The younger age cohorts did report that they preferred an easy-to-use interface, where the interactive device did not detract from the user experience. We coded 57 responses from the younger age cohorts that mentioned interface functionality, in both positive and negative light. As described here, “it would have been better if you did not have to scroll down” (28-43 age cohort, Design One), or it “felt like a bit of a gimmick, why not just have the information under the pictures instead of needing to click nine times” (18-27, Design Three). The data suggests that regardless of the audiences’ age, usability influences engagement. For the older age cohorts, however, the effects of poor usability cause greater effect than a minor aggravation, but feelings of confusion and distraction.
@@
In older audiences, it has been shown that complex visualizations can be cognitively demanding, requiring users to remember and interpret multiple information pieces simultaneously.1 We found evidence that older audiences expressed a preference for less complex visualization, for example, “I don’t find it particularly easy to interpret graphs or charts and I loathe Venn diagrams, so, I possibly had to concentrate more than other participants in order to ensure that I was interpreting the information correctly” (60+ age cohort, Design Two) or “I found the charts a little confusing to begin with - possibly my age!” (44-59 age cohort, Design One). Furthermore, older audiences mentioned that they were required to revisit the narrative visualization to fully comprehend it. We coded 9 instances where participants from the older age cohorts mentioned they missed data in the interactive visualization, this was compared one instance that appeared in the 28-43 age cohort and none were observable in the 18-27 age cohort. “My biggest problem with it was the need to scroll down. At first I didn’t realize there was more data below” (60+ age cohort, Design One). This reveals older audiences might miss crucial information if the representation is too complex or the interactive device is not clearly marked.
@@
Qualitative data analysis indicated a stark difference in the terminology used by the youngest age cohort relative to the older age cohorts. Interactivity was mentioned 49 times by the youngest age cohort, and usually in a positive light. Furthermore, the youngest age cohort seemed to have a perceptive understanding of how the interactive narrative pattern achieved its intent of encouraging engagement. For example, the younger age cohort’s responses to Design Three, ‘Exploration’ specifically outlined how exploring the data through clicking was more engaging than simply reading it. Our study sheds light on the depth of understanding of interactive devices held by the youngest age cohort, illustrated by their descriptive feedback. The results of this study do not disprove the NN Group report, which stated it was a myth that young adults “crave multimedia and innovative design.”[20] Rather, the results of this study give credence to the fact that young adults are accustomed and therefore more understanding of interactive devices. Increased engagement is, therefore largely due to the young adults’ ability to perceive the intent of the author. Whereby recognizing that as an audience, younger age cohorts are expected to be engaged therefore they are engaged. For future authors of interactive narrative visualization targeted towards younger age cohorts, it is recommended to use interactive narrative patterns with explicit intent. Duplicitous or superfluous use of interactive narrative patterns would likely be recognized and thus could result in lowered engagement.

### v2

@@ Abstract
To achieve our aim, we performed an empirical experiment. We developed three narrative visualization examples that employed different interactive narrative patterns tailored specifically for engagement [8]. We randomly assigned 2400 participants to one of the three narrative visualization examples and measured their engagement using VisEngage [6]. VisEngage is a self-reporting questionnaire specifically developed to measure engagement in visualization. The age groups were split into generations. A younger audience with an age of 18-27, a middle-younger audience between the ages of 28-43, an older audience of 44-59, and finally, the oldest age cohort consisting of over 60. These age groups were determined by generation boundaries.
@@
Previous work on different age cohorts in visualization usually focused on either the very old or the very young. For example, visualization research with children has investigated pedagogical approaches for visual literacy and visualization design. Visualization research into elderly audiences has examined aspects of accessibility, comprehension and perception. The lack of understanding of the needs of different age cohorts is a known and cited dilemma in visualization research [2].
In the HCI field, age-based research has found significant differences between age groups. For example, the time taken to perform input modalities of end-users compared to their age group found that older adults were significantly different from their younger counterparts. Although the results were inconclusive, strong evidence pointed to differences between older and younger age cohorts when using devices such as smart watches [18].
@@
We conducted a crowd-sourced study using three narrative visualization examples as a stimulus to achieve our research aim of investigating if and how, end-user age impacts engagement in interactive narrative visualization. To see the interactive narrative visualization example code and raw data, please see the supplementary material here.
@@
The second narrative visualization example, Design Two, used the ‘Breaking the Fourth Wall’ interactive narrative pattern. ‘Breaking the Fourth Wall’ is a term often cited in cinema and literature disciplines. In interactive narrative visualization, a direct question is asked of the audience, normally to input personal data. This creates a ‘self-story connection,’ which has been found to encourage engagement, as it includes the user within the story [4]. The narrative visualization design example that inspired this study was a finalist in the online journalism awards. Published by the ABC Australia Story Lab, the narrative visualization is titled ‘See how global warming has changed the world since your childhood.’ In our study, Design Two broke the fourth wall by asking the user to input their name. Specifically, the user was asked to ‘please enter your first name (this data will not be stored)’ so that privacy concerns were availed with the assurance that data relating to the user’s name would not be stored. Design Two opened by stating, ‘How likely are you to be murdered by a serial killer?’ The user was asked to enter their first name click submit. Once submitted, a screen appeared with ‘[name], have you ever watched a horror movie and found yourself scared? Should you be scared of serial killers?’ This is followed by a basic bar chart explaining that ’1% of homicides in the UK are about 5.’ The narrative sequentially revealed itself as the user scrolled down the screen.’ All data sources are referenced at the end of the narrative visualization. See Figure 1:B for a diagram of Design Two.
@@ Results
‘Interactive’ was explicitly mentioned 7 times by the 60+ cohort. This was in contrast to the 18-27 age cohort, who explicitly mentioned ‘interactive’ or ‘interactivity’ 49 times. We analyzed the exact phrases that participants used in the 18-27 cohort relating to interactivity. Our analysis revealed that the younger audience attributed their engagement to interactivity, for example, “yes because it was interactive” or I was engaged as it was an interactive task” (both comments from 18-27 age cohort, Design One). The interactive device in Design Two, aimed to include the audience in the story and thus encourage engagement. We coded 22 instances where the 18-24 age cohort recognized that this was the aim of the interactive device in Design 2. For example, “Yes as it was interactive and by using my name felt personal” (18-27 age cohort, Design Two). In contrast, we coded 3 instances in the 60+ age cohort specifically mentioning the interactive device in Design Two. We found further evidence of perceptiveness in Design Three from the 18-27 age cohort, where we coded 18 responses that noticed the interactive device, which in this case required clicking and exploring the interface. For example, “I did feel I was engaged in the data story as I had to click to find the information as well as scroll for more information” (18-27 age cohort, Design Three). The goal of the interactive device in Design One, ‘Make a Guess’ was to illustrate a mismatch between audience expectation and reality. We coded 17 responses from the 18-27 age cohort that referenced the interactive device. For example, ‘The decline is higher’ which made me feel engaged” (18-27 age cohort, Design One). In contrast the 60+ age cohort mentioned the interactive device in Design One 6 times. While we observed that 18-24 age cohort was relatively more aware of the interactive devices this does not mean other age groups were oblivious, only their perceptive feedback was less frequent. For example, “I felt engaged as there was an interactive question where I could enter what I thought to be the answer. This made the impact of learning the true answer heavier as I was engaging with the story” (44-59 age cohort, Design One) or “I found the personality types interesting flip over to read and to associate the descriptions with the images you gave for the personality” (60+ age cohort, Design Three). When recognized by the older age cohorts, there were 17 instances where the interactive device had the opposite effect of encouraging engagement and reported as a distraction or confusion, “I was distracted by, initially, not realizing I had to scroll down the box to gain more information” (60+ age cohort, Design Two) or “I found the chart near the end distracting as it seemed confusing.” (44-59 age cohort, Design Three). In the 18-27 age cohort, there were 2 observable reports of the interactive devices causing confusion and none causing distraction. The two reports were varied on the reasoning for why the interactive narrative visualization was deemed confusing. The younger age cohorts did report that they preferred an easy-to-use interface, where the interactive device did not detract from the user experience. We coded 57 responses from the younger age cohorts that mentioned interface functionality, in both positive and negative light. As described here, “it would have been better if you did not have to scroll down” (28-43 age cohort, Design One), or it “felt like a bit of a gimmick, why not just have the information under the pictures instead of needing to click nine times” (18-27, Design Three). The data suggests that regardless of the audiences’ age, usability influences engagement. For the older age cohorts, however, the effects of poor usability cause greater effect than a minor aggravation, but feelings of confusion and distraction.
@@
In older audiences, it has been shown that complex visualizations can be cognitively demanding, requiring users to remember and interpret multiple information pieces simultaneously. We found evidence that older audiences expressed a preference for less complex visualization, for example, “I don’t find it particularly easy to interpret graphs or charts and I loathe Venn diagrams, so, I possibly had to concentrate more than other participants in order to ensure that I was interpreting the information correctly” (60+ age cohort, Design Two) or “I found the charts a little confusing to begin with - possibly my age!” (44-59 age cohort, Design One). Furthermore, older audiences mentioned that they were required to revisit the narrative visualization to fully comprehend it. We coded 9 instances where participants from the older age cohorts mentioned they missed data in the interactive visualization, this was compared one instance that appeared in the 28-43 age cohort and none were observable in the 18-27 age cohort. “My biggest problem with it was the need to scroll down. At first I didn’t realize there was more data below” (60+ age cohort, Design One). This reveals older audiences might miss crucial information if the representation is too complex or the interactive device is not clearly marked.
@@
Qualitative data analysis indicated a stark difference in the terminology used by the youngest age cohort relative to the older age cohorts. Interactivity was mentioned 49 times by the youngest age cohort, and usually in a positive light. Furthermore, the youngest age cohort seemed to have a perceptive understanding of how the interactive narrative pattern achieved its intent of encouraging engagement. For example, the younger age cohort’s responses to Design Three, ‘Exploration’ specifically outlined how exploring the data through clicking was more engaging than simply reading it. Our study sheds light on the depth of understanding of interactive devices held by the youngest age cohort, illustrated by their descriptive feedback. The results of this study do not disprove the NN Group report, which stated it was a myth that young adults “crave multimedia and innovative design.”[20] Rather, the results of this study give credence to the fact that young adults are accustomed and therefore more understanding of interactive devices. Increased engagement is, therefore largely due to the young adults’ ability to perceive the intent of the author. Whereby recognizing that as an audience, younger age cohorts are expected to be engaged, therefore they are engaged. For future authors of interactive narrative visualization targeted towards younger age cohorts, it is recommended to use interactive narrative patterns with explicit intent. Duplicitous or superfluous use of interactive narrative patterns would likely be recognized and thus could result in lowered engagement.